# José Luis Blanco-Claraco

### v1

@@ 2.4 LiDAR-based loop closure and SLAM
What does actually mean a loop closure in practice, and the way the problem is approached, are both strongly influenced by the employed sensors. For visual SLAM with typical pin-hole cameras, moving a camera around a small room would lead to loop closures, while for a 3D LiDAR with 360 of horizontal field of view one might never consider a loop closure situation if moving in a large warehouse if there are no severe occlusions. This variety of situations makes this research field extremely active for each sensor niche (Huang (2019); Wang et al. (2019); Arshad and Kim (2021); Tsintotas et al. (2022)).
Regarding the topic of this work, LiDAR sensors, a recent work by Kim et al. (2021) introduced a new rotation-invariant descriptor for performing topological and (partial) metric localization. Similar ideas are proposed in related works like LiDAR-IRIS (Wang et al. (2020)) which uses a polar representation of 3D point clouds, or in Gupta et al. (2024).
@@ 3 Proposed architecture
This section describes in detail the different parts of the proposed system. The next subsection starts providing a top-level overview of the key elements, and subsequently they are explained individually from a scientific and algorithmic point of view. For a software engineering overview, refer to Appendix A and to the online documentation linked therein.
@@ 3.1 Overview
Pursuing the goals defined in the Introduction, we define the basic blocks illustrated in Figure 1.
Among them, three essential types of procedural blocks or algorithms are defined. Firstly, we have the odometry and loop-closure subsystems. The idea of splitting SLAM into odometry (“local mapping”) and loop-closure (“global mapping”) can be traced back to the visual SLAM community with PTAM (Klein and Murray (2007)) and has been adopted by most subsequent successful solutions for visual SLAM (RTAB-Map Labbe and Michaud (2014), ORB-SLAM Campos et al. (2021)) or LiDAR SLAM (Cartographer Hess et al. (2016)). These two blocks are discussed in §3.5 and §3.13, respectively.
Secondly, we define generic metric mapping pipelines as one of the core elements in our framework, since they are used internally as building blocks for both, odometry and loop-closure, apart of having additional applications. They can be thought of as arbitrary combinations of algorithm as processing nodes in a data-driven network, with data being metric maps. Their final goal is taking raw sensory data (or a former metric map) as inputs, and processing them to create a new metric map (or to modify the existing one). They are introduced in §3.4.
@@
: This represents the source of sensor data, which can be either a real live device or an offline dataset. Real-world applications will normally use the former, while benchmarking and development predominantly use the latter for convenience. Available options in our framework are enumerated in §3.2
@@
: Most existing methods in the literature use one single metric map representation (e.g. point clouds in Deschaud (2018) or Vizzo et al. (2023)) while others maintain two (e.g. corner and plane points in LOAM (Zhang and Singh (2017)) and most derived works). The present work proposes allowing the definition of an arbitrary set of metric map representations and data structure implementations. Some of them will be more useful during real-time mapping, others during map postprocessing or loop closing, others for obstacle avoidance, etc. More on metric maps follows up in §3.3
@@
: Dubbed “simple-maps” in our implementation, they contain a subset of all incoming raw sensor data, synchronized and paired with the estimated pose according to an odometry or SLAM method to form key-frames. More details are given in §3.10.
@@
where, following the RIGID notation for poses (Nadeau (2024)), $\hskip 0.80002pt\prescript{}{w}{\boldsymbol{T}}\prescript{}{i}{}\equiv\boldsymbol{T}\prescript{}{i}{}$ stands for the $i$-th SE(3) pose of the vehicle local reference frame with respect to the world ($w$), that is, in global coordinates.
3.2 Dataset sources
The former work in Blanco (2019) already introduced data source MOLA modules as an abstraction of either, real live sensors or offline datasets. Such abstraction aims at enabling algorithm implementations in MOLA to immediately work with any of the supported datasets, without worrying about the details of each dataset file and data structures, sensor intrinsic and extrinsic calibration, ground truth (GT) axes conventions, etc.
  • MRPT rawlog files: Rawlog files are an equivalent to ROS bag files, introduced in the MRPT framework in 2005. It has been used in public datasets, e.g. Blanco-Claraco et al. (2014, 2019). Backwards compatibility and Operative System (OS) independence has been ensured during the whole life of the project, hence software written today using the latest MRPT version is able to read all past datasets, even if generated in a different OS or processor architecture. There exists an import tool from ROS 1 bag files to rawlogs, hence this MOLA input module also offers easy access to many of the ROS 1 public datasets, e.g. HILTI challenge datasets (Helmberger et al. (2022)) or the Newer College dataset (Zhang et al. (2021)).
Finally, a word is in order regarding memory management for dataset input modules. Loading and keeping in memory whole datasets may become unfeasible for a typical desktop computer, e.g. some individual sequences of the MulRan dataset (Kim et al. (2020)) take 29 GiB. The present framework exploits the usage of MRPT lazy-load mechanism in different areas, including dataset sources. Observations that tend to be heavy in memory requirements (i.e. images, 3D LiDAR scans, and RGB+D frames) can be automatically loaded to memory when accessed, to be unloaded once a fixed number of additional frames have been processed. Most MOLA dataset input modules also feature read-ahead in a multi-thread fashion to minimize the performance impact of such lazy-load mechanism. In practice, this creates a cache of recently-used and about-to-be-used sensory data quickly accessible in RAM while SLAM is processing them.
@@ 3.3 Metric maps
The idea is using such compound metric maps as the fundamental data containers used as both inputs and outputs of metric map pipelines (see §3.4) and as inputs to the ICP-like optimization pipeline (see §3.7). The existence of different map layers is justified by two facts:
@@
Such layer differentiation is useful during mapping, but also in posterior post-processing stages. Our framework allows the definition of new metric map types by users by means of a virtual base interface that abstracts all common operations of maps such as clearing, inserting sensor observations, evaluating the likelihood of a given observation, searching for nearest neighbors, etc. Having such abstract interface does not only allow for the extensibility and generic programming in our framework, but also enables performing performance benchmarking to take optimal decisions about what map type to use for each application. Next we enumerate the metric map types that have been used in the experiments presented in this work  [^1] The list of C++ class names and corresponding libraries where each map type is defined in Appendix A.:
@@
    Point clouds with contiguous storage layout in memory for each point component: This group of types include maps for simple clouds where each point only has $(X,Y,Z)$ components, with an intensity component $(X,Y,Z,I)$, and with LiDAR ring identifier and per-point timestamps $(X,Y,Z,I,R,T)$.
@@
    Point clouds with a hashed voxels storage layout: this map type implements 3D voxels indexed by voxel coordinates using the optimal hash function introduced in Teschner et al. (2003), with a hard-coded maximum number of points. This is the most common map layer type in most parts of the proposed system, and it is based on similar successful implementations in past works (Deschaud (2018); Vizzo et al. (2023)).
@@ 3.4 Metric map generation and update pipelines
Metric map pipelines are at the core of the flexibility offered by the present framework. Such pipelines are defined as a bipartite, directed graphs with two kinds of nodes: data nodes (metric maps, as defined in §3.3) and action nodes (either filters or generators). Directed edges may only exist between data nodes and action nodes. Follow Figure 2 as a reference for the discussion below, keeping in mind that despite map layers are represented there as point clouds, arbitrary map types are allowed as long as they are compatible with their action nodes.
Figure 2: Metric map pipelines represent map processing networks as bipartite graphs with two kinds of nodes: data (map layers) and actions (generators or filters). Different connection patterns exist: (a) generators converting raw observations into map layers, and filters of different arity such as (b) unary, (c) binary, or (d) ternary map filters. See discussion in §3.4.
Generators (see Figure 2(a)) are especial actions as they are the entry points for sensor data, transforming raw range data, images, depth images, etc. into more elaborated data structures, e.g. point clouds with different per-point field annotations, or plane patches normals. They can also generate more than one output metric map layers, becoming the ideal location where to run classification tasks, e.g. detecting edges or special objects in depth cameras or LiDAR observations. We provide hooks for implementing custom detection algorithms that need to exploit the particular data structure for each possible input sensor, e.g. edge detection algorithms from LiDAR range images. A generic default generator exists which just creates a map layer of a given type (by default, a point cloud) and inserts the observation into it by using the abstract interface mentioned in §3.3. For 2D or 3D LiDARs, this simply creates an unfiltered point cloud with all sensed points in a given scan.
Filters are the other kind of action in this framework, and are broadly defined as any arbitrary algorithm that takes at least one map layer as input and generates or modifies other (or the same input) layers. Our work already ships several such filters, while also offering a plug-in system to add user-defined ones. Some of the most important filters, used in the LiDAR odometry pipeline presented in §3.8, are briefly introduced next:
@@ 3.4.1 De-skewing
with ${\cdot}^{\wedge}$ the hat or wedge operator, mapping SO(3) tangent space vectors to rotation matrices (Blanco-Claraco (2021)).
@@ 3.4.3 Down-sampling
: As shown while discussing related LO works (§2.1), sub-sampling point clouds is essential to achieve stable ICP-based mapping. Our framework offers several down-sampling filters, ranging from simply picking the first point within each 3D voxel for a given uniform volumetric resolution, to keeping the average of each such voxels, or using non-linear spatial tessellation.
@@ 3.5 LiDAR odometry module
As discussed earlier, this work follows a scan-to-map design pattern since it has proven benefits for 3D LiDAR mapping. Hence, after using raw observations to build a metric map (block #7), it is aligned against a local metric map (block #9), which is initialized via a special instance of a metric map pipeline (block #5) that normally will just create empty metric map layers. This block is run only once at system start up or when it is reset. The local metric map data structure (block #9), is selectively updated using a custom pipeline algorithm (block #8) using one or more layers from the observation metric map (block #7) when certain criteria are met, as determined by a so-called local map update decider (block #14). An independent decider (block #15), normally with different criteria, is used to determine what frames become key-frames in the output simple-map. In general, simple-maps would benefit from sparser key-frames to reduce the size and memory requirements of final maps, while the local map tends to benefit from more frequent updates from incoming sensor frames. However, note that most past works (e.g. Deschaud (2018); Vizzo et al. (2023)) update the local map for every single incoming sensor frame. This may be optimal in driving scenarios, but our experimental results show that this policy may make the system to diverge when using 3D LiDARs indoors, specially in tight spaces.
Pose tracking itself is achieved by means of an ICP-like iterative optimizer (blocks #11 and #12), in charge of finding the SE(3) pose $\mathbf{q}[k]$ that best explains the observation given a local map. This optimization procedure is detailed in §3.7. For now, it is enough to abstract it as an algorithm taking as inputs: (i) the initial guess $\mathbf{q}^{\prime}[k]$ from the kinematic state estimator, (ii) a local and a global metric map (with different geometric entities and map layers depending on the application), and (ii) a set of dynamic variables (block #10) to adapt the behavior of the internal ICP pipeline (see §3.6).
Once the optimizer converges, a set of quality evaluation algorithms (block #13; see §3.7.4) are used to estimate the quality of the optimization. This metric, together with other conditions (not shown in the diagram for the sake of clarity) are then used to conditionally trigger the two decider blocks: (i) one to accept the alignment as good enough to append the obtained pose to the estimated trajectory (block #4) and to update the local metric map (block #9) via the purposely designed pipeline (block #8), and (ii) another one (block #15) to add new key-frames to the output simple-map (block #16). Note that a localization-only mode exist (see §6), which disables both deciders while still producing an estimate of the vehicle state for each time step in $\mathbf{q}[k]$ by aligning the current observation against a prebuilt local map.
@@ 3.6 Adaptive behavior: dynamic variables
All software blocks in the LO system (Figure 3) have parameters that are initialized at system start-up. Some such parameters, however, would benefit from being dynamically adapted depending on sensed conditions like the size of the environment (indoors vs. outdoors) or the motion profile (driving, handheld, drone, etc.). For example, take the voxel-based down-sampling mechanism used in IMLS-SLAM (Deschaud (2018)), which used a fixed resolution of 1.0 meters, adequate for driving datasets. If the SLAM system is initialized in a small office-like environment instead, a much smaller resolution (i.e. less down-sampling) would be needed indeed. Instead of relying on manual tuning of all those parameters, our system would automatically detect the approximate size of the environment and initialize the local map and down-sampling resolutions accordingly.
To make this adaptation possible, our framework offers two features: (i) supporting mathematical expressions and basic scripting directly in the parameter specification files, and (ii) defining dynamic variables that can be used in such expressions. On the former, we use the C++ Mathematical Expression Toolkit Library (ExprTk) (Partow (2015)) due to its lightweight and efficient implementation. On the latter, a set of variables are defined so they can be used as symbols in the expressions, which are re-evaluated for each time step, and even for each ICP iteration within a given time step. Next we review the most relevant such variables and how they are updated.
@@ 3.6.3 Adaptive matching threshold
: We have incorporated the idea of a dynamic threshold, proposed in KISS-ICP, to parameterize the data association and optimization stages of the ICP optimization loop according to perceived changes. As justified geometrically in Vizzo et al. (2023), given translational and rotational corrections $\boldsymbol{\Delta}\mathbf{p}\in\mathbb{R}^{3}$ and $\boldsymbol{\Delta}\mathbf{R}\in SO(3)$, respectively, of the ICP optimized pose with respect to the initial guess from the kinematic state prediction (§3.9), the expected maximum point-to-point error $\delta$ can be approximated with:
$\delta=\left\lVert\boldsymbol{\Delta}\mathbf{p}\right\rVert+2r_{max}\sin\frac{1}{2}\left\lVert\log(\boldsymbol{\Delta}\mathbf{R})^{\vee}\right\rVert$ (4)
where $(\cdot)^{\vee}$ is the vee-operator, the inverse of the wedge operator used in Eq. (2) , $\log\mathbf{R}$ is the matrix logarithm, and the averaged maximum sensor range $r_{max}$ is described in §3.6.1. With the idea of obtaining a threshold $\sigma$ to scale data association and optimization within ICP out of sequences of $\delta$ values over time, KISS-ICP proposes taking the standard deviation of all $\delta$ values, saturating with a minimum $\sigma_{min}$ to prevent degeneration after long runs with good motion model predictions. While quite successful for driving scenarios, this approach lacks the ability to quickly react against abrupt changes in conditions (e.g. large occlusions, crossing a door moving from outdoors to an indoor area or handheld sensors with high dynamics). In order to further improve the adaptability of the whole system, we propose using a closed-loop proportional negative feedback controller to cope with sudden perturbations in the quality of ICP alignments, as sketched in Figure 4. This controller generates a multiplicative term $c$ that makes $\delta$ to grow ($\sigma^{\prime}=c\delta$) whenever ICP quality starts to drop, promoting the search for additional pairings in the next time step by increasing the search radius for nearest neighbors. A first-order low pass filter, identical to Eq. (3), is then applied to obtain the actual threshold $\sigma$. As it is well-known in control theory, a proportional controller always exhibits an offset error, hence the obtained ICP quality will never reach the set-point, but in practice most often remains in the range $[0.9,1]$.

### v2

@@ 2.4 LiDAR-based loop closure and SLAM
What does actually mean a loop closure in practice, and the way the problem is approached, are both strongly influenced by the employed sensors. For visual SLAM with typical pin-hole cameras, moving a camera around a small room would lead to loop closures, while for a 3D LiDAR with 360 of horizontal field of view one might never consider a loop closure situation if moving in a large warehouse if there are no severe occlusions. This variety of situations makes this research field extremely active for each sensor niche (Huang, 2019; Wang et al., 2019; Arshad and Kim, 2021; Tsintotas et al., 2022).
Regarding the topic of this work, LiDAR sensors, a recent work by Kim et al. (2021) introduced a new rotation-invariant descriptor for performing topological and (partial) metric localization. Similar ideas are proposed in related works like LiDAR-IRIS (Wang et al., 2020) which uses a polar representation of 3D point clouds, or in Gupta et al. (2024).
@@ 3 Proposed architecture
This section describes in detail the different parts of the proposed system. The next subsection starts providing a top-level overview of the key elements, and subsequently they are explained individually from a scientific and algorithmic point of view.
@@ 3.1 Overview
Pursuing the goals defined in the Introduction, we define the basic blocks illustrated in Figure 1. Among them, three essential types of procedural blocks or algorithms are defined. Firstly, we have the odometry and loop-closure subsystems. The idea of splitting SLAM into odometry (“local mapping”) and loop-closure (“global mapping”) can be traced back to the visual SLAM community with PTAM (Klein and Murray, 2007) and has been adopted by most subsequent successful solutions for visual SLAM (RTAB-Map Labbe and Michaud (2014), ORB-SLAM Campos et al. (2021)) or LiDAR SLAM (Cartographer Hess et al. (2016)). These two blocks are discussed in Section 3.4 and Section 3.12, respectively.
Secondly, we define generic metric mapping pipelines as one of the core elements in our framework, since they are used internally as building blocks for both, odometry and loop-closure, apart of having additional applications. They can be thought of as arbitrary combinations of algorithm as processing nodes in a data-driven network, with data being metric maps. Their final goal is taking raw sensory data (or a former metric map) as inputs, and processing them to create a new metric map (or to modify the existing one). They are introduced in Section 3.3.
@@
: This represents the source of sensor data, which can be either a real live device or an offline dataset. Real-world applications will normally use the former, while benchmarking and development predominantly use the latter for convenience. Available options in our framework are enumerated in Section 3.13
@@
: Most existing methods in the literature use one single metric map representation (e.g. point clouds in Deschaud (2018) or Vizzo et al. (2023)) while others maintain two (e.g. corner and plane points in LOAM (Zhang and Singh, 2017) and most derived works). The present work proposes allowing the definition of an arbitrary set of metric map representations and data structure implementations. Some of them will be more useful during real-time mapping, others during map postprocessing or loop closing, others for obstacle avoidance, etc. More on metric maps follows up in Section 3.2
@@
: Dubbed “simple-maps” in our implementation, they contain a subset of all incoming raw sensor data, synchronized and paired with the estimated pose according to an odometry or SLAM method to form key-frames. More details are given in Section 3.9.
@@
where, following the RIGID notation for poses (Nadeau, 2024), $\hskip 0.80002pt\prescript{}{w}{\boldsymbol{T}}\prescript{}{i}{}\equiv\boldsymbol{T}\prescript{}{i}{}$ stands for the $i$-th SE(3) pose of the vehicle local reference frame with respect to the world ($w$), that is, in global coordinates.
@@ 3.2 Metric maps
The idea is using such compound metric maps as the fundamental data containers used as both inputs and outputs of metric map pipelines (see Section 3.3) and as inputs to the ICP-like optimization pipeline (see Section 3.6). The existence of different map layers is justified by two facts:
@@
Such layer differentiation is useful during mapping, but also in posterior post-processing stages. Our framework allows the definition of new metric map types by users by means of a virtual base interface that abstracts all common operations of maps such as clearing, inserting sensor observations, evaluating the likelihood of a given observation, searching for nearest neighbors, etc. Having such abstract interface does not only allow for the extensibility and generic programming in our framework, but also enables performing performance benchmarking to take optimal decisions about what map type to use for each application. Next we enumerate the metric map types that have been used in the experiments presented in this work:
@@
    Point clouds with contiguous storage layout in memory for each point component: This group of types include maps for simple clouds where each point only has $(X,Y,Z)$ components, with an intensity ($I$) component $(X,Y,Z,I)$, and with LiDAR ring ($R$) identifier and per-point timestamps ($T$), that is, $(X,Y,Z,I,R,T)$ components.
@@
    Point clouds with a hashed voxels storage layout: this map type implements 3D voxels indexed by voxel coordinates using the optimal hash function introduced in Teschner et al. (2003), with a hard-coded maximum number of points. This is the most common map layer type in most parts of the proposed system, and it is based on similar successful implementations in past works (Deschaud, 2018; Vizzo et al., 2023).
@@
    3D-NDT (Normal Distribution Transform) maps, as proposed in Magnusson et al. (2007), where a 3D sparse voxel map stores fitted Gaussian distributions for points in each voxel. In our implementation, voxels with a sufficient number of points whose spatial distribution clearly defines a plane store the mean and covariance matrix of such points, whereas voxels with too few samples or which are not planar enough retain all individual points. The idea is to enable our map to establish both, point-to-point pairings (for voxels which are not planar) and point-to-plane pairings (for planar ones). In particular, the NDT representation for a voxel is used if $\sigma_{1}/\sigma_{2}<\tau$, with $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ the three eigenvalues of the voxel point distribution sorted in ascending order and $\tau$ a threshold (0.05 in our experiments). As the map is populated with incoming observations, individual voxels may switch between the two representations.
  5. 5.
@@ 3.3 Metric map generation and update pipelines
Metric map pipelines are at the core of the flexibility offered by the present framework. Such pipelines are defined as a bipartite, directed graphs with two kinds of nodes: data nodes (metric maps, as defined in Section 3.2) and action nodes (either filters or generators). Directed edges may only exist between data nodes and action nodes. Follow Figure 2 as a reference for the discussion below, keeping in mind that despite map layers are represented there as point clouds, arbitrary map types are allowed as long as they are compatible with their action nodes.
Figure 2: Metric map pipelines represent map processing networks as bipartite graphs with two kinds of nodes: data (map layers) and actions (generators or filters). Different connection patterns exist: (a) generators converting raw observations into map layers, and filters of different arity such as (b) unary, (c) binary, or (d) ternary map filters. See discussion in Section 3.3.
Generators (see Figure 2(a)) are especial actions as they are the entry points for sensor data, transforming raw range data, images, depth images, etc. into more elaborated data structures, e.g. point clouds with different per-point field annotations, or plane patches normals. They can also generate more than one output metric map layers, becoming the ideal location where to run classification tasks, e.g. detecting edges or special objects in depth cameras or LiDAR observations. We provide hooks for implementing custom detection algorithms that need to exploit the particular data structure for each possible input sensor, e.g. edge detection algorithms from LiDAR range images. A generic default generator exists which just creates a map layer of a given type (by default, a point cloud) and inserts the observation into it by using the abstract interface mentioned in Section 3.2. For 2D or 3D LiDARs, this simply creates an unfiltered point cloud with all sensed points in a given scan.
Filters are the other kind of action in this framework, and are broadly defined as any arbitrary algorithm that takes at least one map layer as input and generates or modifies other (or the same input) layers. Our work already ships several such filters, while also offering a plug-in system to add user-defined ones. Some of the most important filters, used in the LiDAR odometry pipeline presented in Section 3.7, are briefly introduced next:
@@ 3.3.1 De-skewing
with ${\cdot}^{\wedge}$ the hat or wedge operator, mapping SO(3) tangent space vectors to rotation matrices (Blanco-Claraco, 2021).
@@ 3.3.3 Down-sampling
: As shown while discussing related LO works (Section 2.1), sub-sampling point clouds is essential to achieve stable ICP-based mapping. Our framework offers several down-sampling filters, ranging from simply picking the first point within each 3D voxel for a given uniform volumetric resolution, to keeping the average of each such voxels, or using non-linear spatial tessellation.
@@ 3.4 LiDAR odometry module
As discussed earlier, this work follows a scan-to-map design pattern since it has proven benefits for 3D LiDAR mapping. Hence, after using raw observations to build a metric map (block #7), it is aligned against a local metric map (block #9), which is initialized via a special instance of a metric map pipeline (block #5) that normally will just create empty metric map layers. This block is run only once at system start up or when it is reset. The local metric map data structure (block #9), is selectively updated using a custom pipeline algorithm (block #8) using one or more layers from the observation metric map (block #7) when certain criteria are met, as determined by a so-called local map update decider (block #14). An independent decider (block #15), normally with different criteria, is used to determine what frames become key-frames in the output view-based map. In general, such maps would benefit from sparser key-frames to reduce the size and memory requirements of final maps, while the local map tends to benefit from more frequent updates from incoming sensor frames. However, note that most past works (e.g. Deschaud (2018); Vizzo et al. (2023)) update the local map for every single incoming sensor frame. This may be optimal in driving scenarios, but our experimental results show that this policy may make the system in other scenarios, as analyzed in Section 7.2.
Pose tracking itself is achieved by means of an ICP-like iterative optimizer (blocks #11 and #12), in charge of finding the SE(3) pose $\mathbf{q}[k]$ that best explains the observation given a local map. This optimization procedure is detailed in Section 3.6. For now, it is enough to abstract it as an algorithm taking as inputs: (i) the initial guess $\mathbf{q}^{\prime}[k]$ from the kinematic state estimator, (ii) a local and a global metric map (with different geometric entities and map layers depending on the application), and (ii) a set of dynamic variables (block #10) to adapt the behavior of the internal ICP pipeline (see Section 3.5).
Once the optimizer converges, a set of quality evaluation algorithms (block #13; see Section 3.6.4) are used to estimate the quality of the optimization. This metric, together with other conditions (not shown in the diagram for the sake of clarity) are then used to conditionally trigger the two decider blocks: (i) one to accept the alignment as good enough to append the obtained pose to the estimated trajectory (block #4) and to update the local metric map (block #9) via the purposely designed pipeline (block #8), and (ii) another one (block #15) to add new key-frames to the output map (block #16). Note that a localization-only mode exist (see Section 6), which disables both deciders while still producing an estimate of the vehicle state for each time step in $\mathbf{q}[k]$ by aligning the current observation against a prebuilt local map.
@@ 3.5 Adaptive behavior: dynamic variables
All software blocks in the LO system (Figure 3) have parameters that are initialized at system start-up. Some such parameters, however, would benefit from being dynamically adapted depending on sensed conditions like the size of the environment (indoors vs. outdoors) or the motion profile (driving, handheld, drone, etc.). For example, take the voxel-based down-sampling mechanism used in IMLS-SLAM (Deschaud, 2018), which used a fixed resolution of 1.0 meters, adequate for driving datasets. If the SLAM system is initialized in a small office-like environment instead, a much smaller resolution (i.e. less down-sampling) would be needed indeed. Instead of relying on manual tuning of all those parameters, our system would automatically detect the approximate size of the environment and initialize the local map and down-sampling resolutions accordingly.
To make this adaptation possible, our framework offers two features: (i) supporting mathematical expressions and basic scripting directly in the parameter specification files, and (ii) defining dynamic variables that can be used in such expressions. On the former, we use the C++ Mathematical Expression Toolkit Library (ExprTk) (Partow, 2015) due to its lightweight and efficient implementation. On the latter, a set of variables are defined so they can be used as symbols in the expressions, which are re-evaluated for each time step, and even for each ICP iteration within a given time step. Next we review the most relevant such variables and how they are updated.
@@ 3.5.3 Adaptive matching threshold
: We have incorporated the idea of a dynamic threshold, proposed in KISS-ICP, to parameterize the data association and optimization stages of the ICP optimization loop according to perceived changes. As justified geometrically in Vizzo et al. (2023), given translational and rotational corrections $\boldsymbol{\Delta}\mathbf{p}\in\mathbb{R}^{3}$ and $\boldsymbol{\Delta}\mathbf{R}\in SO(3)$, respectively, of the ICP optimized pose with respect to the initial guess from the kinematic state prediction (Section 3.8), the expected maximum point-to-point error $\delta$ can be approximated with:
$\delta=\left\lVert\boldsymbol{\Delta}\mathbf{p}\right\rVert+2r_{max}\sin\frac{1}{2}\left\lVert(\log\boldsymbol{\Delta}\mathbf{R})^{\vee}\right\rVert$ (4)
where $\log\mathbf{R}$ is the matrix logarithm, the averaged maximum sensor range $r_{max}$ is described in Section 3.5.1, and $(\cdot)^{\vee}$ is the vee-operator, the inverse of the wedge operator used in Eq. (2); that is, the $3\times 3$ rotation correction matrix $\boldsymbol{\Delta}\mathbf{R}$ is first mapped to the tangent space via the matrix logarithm giving a $3\times 3$ skew-symmetric matrix with only three degrees of freedom, then the vee-operator simply takes out those three values as a $3\times 1$ vector.
With the idea of obtaining a threshold $\sigma$ to scale data association and optimization within ICP out of sequences of $\delta$ values over time, KISS-ICP proposes taking the standard deviation of all $\delta$ values, saturating with a minimum $\sigma_{min}$ to prevent degeneration after long runs with good motion model predictions. While quite successful for driving scenarios, this approach lacks the ability to quickly react against abrupt changes in conditions (e.g. large occlusions, crossing a door moving from outdoors to an indoor area or handheld sensors with high dynamics). In order to further improve the adaptability of the whole system, we propose using a closed-loop proportional negative feedback controller to cope with sudden perturbations in the quality of ICP alignments, as sketched in Figure 4. This controller generates a multiplicative term $c$ that makes $\delta$ to grow ($\sigma^{\prime}=c\delta$) whenever ICP quality starts to drop, promoting the search for additional pairings in the next time step by increasing the search radius for nearest neighbors. A first-order low pass filter, identical to Eq. (3), is then applied to obtain the actual threshold $\sigma$. As it is well-known in control theory, a proportional controller always exhibits an offset error, hence the obtained ICP quality will never reach the set-point, but in practice most often remains in the range $[0.9,1]$.